OpenStudy (anonymous):

At one instant, a 17.5-kg sled is moving over a horizontal surface of snow at 3.50 m/s. After 8.75 s has elapsed, the sled stops. Use momentum approach to find the average friction force acting on the sled while it was moving

OpenStudy (anonymous):

\[F \Delta t = m \Delta v\]. then solve for frictional force. F = mv/ t 7N

OpenStudy (anonymous):

note:\(\Sigma F= F_f=\frac{dP}{dT}\)

OpenStudy (anonymous):

where \(P= m\Delta v=m(v-u)\)